# Rise and fall of a multicomponent droplet in a surrounrdfing fluid: simulation study of a bumpy path

### v1

@@ Abstract
The coupling between mass transfer and hydrodynamic phenomena in two-phase flow is not necessarily straightforward due to the different effects that can be encountered. The treatment of such coupling is complex and requires particular efforts, especially in the modelling of the interface between phases. In this paper, we consider the case of a droplet composed of two components (one miscible and one immiscible in water) released in a 2D rectangular domain filled with water. Mass transfer occurs between the miscible element and the surrounding water, which leads to a density inversion that directly affects the droplet trajectory through buoyancy. We perform simulations using a ternary Cahn-Hilliard model (implemented in the “phase_field” model of the TrioCFD code) to capture such coupled phenomena. The Boussinesq approximation for a multicomponent system is used to define the density law and an analytical chemical potential is proposed for the thermodynamic landscape. The effect of the mobility parameter on the flow is highlighted and the results found are in good agreement with the dynamics described from an experimental study of the open literature.
Coupled Cahn-Hilliard/Navier-Stokes equations, Multicomponent System, TrioCFD, Mass Transfer
@@ I Introduction
Multiphase flows are ubiquitous and can be seen in industrial processes and natural environment. They are difficult to describe and model numerically due to the coupling between the moving free surface and the fluid flow. This leads to complex behaviours that are still a challenge to describe and they are therefore driving new development in modelling. This complexity may be augmented when mass transfer between the two phases can occur. This is typically the case in industrial flows and mixing. For instance, such phenomena take place in the corium pool after a severe nuclear accident and affect thermal transfers Zanella et al. (2021); Asmolov et al. (2003); Barrachin and Defoort (2004). Such mass transfer can also occur in microfluidic devices Zhang et al. (2017) and also during the geological sequestration of CO2 either in coal or in deep aquifers De Silva et al. (2015); Zhang and Agarwal (2013). Finally, they are of importance in industrial processes where the dispersion of chemical components in a surrounding fluid is mediated by the motion of droplets Wegener et al. (2014).
@@
In such a situation, mass transfer will occur between the two phases and will affect the fluid flow. This was exemplified in a recent experimental work Rao et al. (2015) inspired by oil spills. Here, through numerical simulations, we show the ability of a ternary phase field model to describe the complex interplay of mass transfer through diffusion and flow equation in the presence of buoyant force. To this purpose, we compare numerical results obtained with a recent experimental work in a qualitative manner. The numerical results also show that the difference between a naive description of the motion and the fully coupled problem can be dramatic and can be well explained by the analysis of the solution of the fully coupled problems. This illustrates the need for fully coupled simulations of such systems. Indeed, while multiphase flows are ubiquitous, simulations methods such as volume of fluids Noh and Woodward (1976) or level set usually rely on the assumption that both phases are immiscible and that no mass transfer should occur through the interface amd S.S. Jain and Dodd (2017). Phase field models based on either the Allen-Cahn Anderson et al. (2000); Folch and Plapp (2005) or the Cahn-Hilliard model Cahn and Hilliard (1958); Kim and Lowengrub (2005); Tóth et al. (2016) can obviously describe mass transfer through the interface. However, this possibility has not been used fully explored and is seen as an adverse effect when modelling multiphase flows of perfectly immiscible fluids Magaletti et al. (2013); Zanella et al. (2020a). Here, we show that a ternary Cahn-Hilliard model coupled with the Navier-Stokes equations, using a simple free energy functional is able to describe the complex interplay between fluid flow and diffusive transport and that such modelling effort brings new insights. This is exemplified by numerical simulations of the motion of a multicomponent droplet moving into water while one of its components diffuses out of the droplet leading to changes in the droplet buoyancy. The numerical results are compared to experimental results presented in Rao et al. (2015).
@@
The paper is organised as follows. First, we briefly describe the phenomena at play in the experiment of interest and discuss qualitatively why a proper description of the coupling between mass transfer and hydrodynamics is needed. Then, we describe the generic form of the model and the simulation setup. Thereafter, we present numerical simulation results that show the ability of the model to describe the interplay between hydrodynamics and mass transfer phenomena before drawing a conclusion and some perspectives.
II Description of the model
@@
Indeed, if the droplet composition is sufficiently rich in the light element, it becomes lighter than water and will move up. However, due to diffusion of the miscible element in water, the amount of A in the droplet will decrease while the amount of heavy element B remains constant. As a result, the density of the droplet will increase until it reaches the density of pure B which is heavier than water. This implies that the density of the droplet that was initially less than water increases, becomes equal to that of the surrounding water and eventually is larger, resulting in the observed motion.
@@
This explanation is fairly simple, however, if one wants to reproduce this behaviour quantitatively, which may be of interest for instance when modelling flow patterns in corium Zanella et al. (2021), the crude approximation where the diffusive mass transfer is computed considering a droplet in pure water and then the motion of the droplet is obviously incorrect, since it neglects the diffusion layer that builds around the droplet. In the same spirit, considering that the mass transfer is the same as for a quiescent droplet is incorrect, since it neglects the advection that affects the composition around the droplet. Therefore, in order to reproduce quantitively such motion, one needs to solve the fully coupled problem: the advection-diffusion equation together with the Navier-Stokes equations. This can be done with numerical simulations using the phase field method Kim and Lowengrub (2005); Tóth et al. (2016); Lee et al. (2021); Vasilopoulos (2020). This class of models relies on a diffuse interface that can be tracked implicitly and has been used to model multiphase flows but rarely to model diffusive transport coupled with flow while it has been shown to be efficient Tree et al. (2017).
@@
It presents the advantage of being thermodynamically consistent: by construction, it derives from a free energy functional (the thermodynamic landscape). In addition, it is fairly simple to implement since it is solely based on coupled PDEs discretised over an Eulerian spatial mesh. Nevertheless, the thermodynamic landscape and model parameters must be chosen with care to give the desired properties. For instance, the diffuse interface thickness is given by a balance between the thermodynamic landscape and the magnitude of a squared gradient term. The accuracy of the model when considering simple multiphase flows has been shown to depend on the mobility of chemical species Magaletti et al. (2013); Zanella et al. (2020a). In the more complex case of ternary system, additional interface properties, such as adsorption, have been shown to be difficult to control Rasolofomanana et al. (2022a) for thermodynamic landscapes obtained from thermodynamic databases such as the ones constructed by the popular Calphad method Cardon et al. (2016a); Rasolofomanana et al. (2022a).
@@
In the present paper, we use the phase field model to study numerically the dynamics of a multicomponent droplet in water. Based on so-called diffuse interface approach, the properties in the phase field model are described by $\phi_{k}$ called “order parameters” which take on two different constant values in each phase and vary continuously within the interface. The diffuse interface description was originally proposed by Van der Waals Rowlinson (1979), it was then popularised by Cahn Cahn and Hilliard (1958) in the context of diffusive mass transfer. It has been then used to describe the motion of free interfaces in a natural way Collins and Levine (1985); Caginalp and Fife (1986) in the context of solid growth where it has become a standard approach Wu (2022); Kadambi et al. (2020). It has been extended thereafter to fluid flow using either a conserved order parameter Anderson et al. (2000) or a non-conserved order parameter Folch and Plapp (2005). Here, since the diffusion of chemical species in a ternary system is considered, we used a conserved approach that allows a simple use of ternary system Tóth et al. (2016) that has already been used in other contexts Tree et al. (2017). In an effort to describe systems of interest, there has been recent efforts to couple the principle of such diffuse interface model with free energy functional extracted from thermodynamic databases Cardon et al. (2016b); Rasolofomanana et al. (2022a). In this paper, we focus on the description of the interplay between diffusive mass transport and fluid flow. Therefore, we have chosen more standard free energy functionals that are prone to very little interface adsorption.
@@
with $\bm{u}$ the advective velocity field, $\tilde{\mu_{j}}$ the chemical potential of the element j, ${G}$ the double-well free energy, the mobility $\mathcal{M}_{i,j}$, the gradient energy $\kappa_{i,j}$ and $\lambda$ an upscaling parameter Cardon et al. (2016b); Rasolofomanana et al. (2022a) defined later. The gradient coefficient matrix is chosen simply to be
@@
where $\kappa^{bin}$ is the gradient energy coefficient obtained from a binary system. The coefficients of the mobility matrix are taken simply to be equal to:
@@
In this work, $M$ is taken to be equal to: $10^{-8}$; $5\times 10^{-9}$, $2.5\times 10^{-9}$, $1\times 10^{-9}$ and $5\times 10^{-10}J^{-1}m^{5}s^{-1}$. We discuss the effect of this parameter on the simulation results in Section IV.
@@
where $P$ is the pressure, $\eta$ is the dynamic viscosity (taken independent of the chemical composition here), $\bm{g}$ is the gravity, the term with $\tilde{\mu_{i}}$ represents the capillary force within the interface as proposed in Kim (2012), $\rho^{\star}$ the mass density of reference and $\rho$ is the mass density that is function of the order parameters $\phi_{A}$ and $\phi_{B}$ and writes, under the assumption of constant molar volume:
@@
with $\rho^{\star}=1000kg/m^{3}$, $\beta_{A}=-0.416$ and $\beta_{B}=0.05$.
@@
where $\phi_{A}^{eq,cont}$ is the composition of the element A in the continuous phase at thermodynamic equilibrium which depends on the size of the domain (see Table 2). With such a choice of thermodynamic landscape, there exists a region where the homogeneous mixture is unstable (the spinodal domain). It corresponds to the domain for which the Hessian matrix of $G$ has one negative eigenvalue and is represented in fig 1. In this region, small fluctuations of the composition can grow and lead to the formation of domains with different compositions. It should be noted that with this choice of ${G}$, no adsorption at the interface between pure B and pure water occurs when using a diagonal $\bar{\bar{\kappa}}$.
$\begin{array}[]{|c|c||c|c|}\hline\cr\eta&10^{-}3\ \mbox{Pa.s}&\sigma&0.0036\ \mbox{J.m}^{-2}\\ \epsilon&0.0008\ \mbox{m}&\lambda&5.4\ \mbox{J.m}^{-3}\\ \rho^{\star}&1000\ \mbox{kg.m}^{-3}&\beta_{A}&-0.416\\ \beta_{B}&0.05&\kappa^{bin}&4.3210^{-5}\ \mbox{J.m}^{-1}\\ \rho_{A}&0.584\ \mbox{kg.m}^{-3}&\rho_{B}&1050\ \mbox{kg.m}^{-3}\\ \hline\cr\end{array}$
@@
The initial compositions of the droplet and of water $\phi_{A}^{ini,dis}$ and $\phi_{B}^{ini,dis}$ must satisfy some conditions. No phase separation should occur during the whole simulation neither in the droplet nor in the bulk phase. Therefore, the initial compositions must be outside the spinodal region and the path between the initial and the equilibrium compositions must not cross it. If ($\phi_{A}^{ini,dis},\phi_{B}^{ini,dis}$) is inside the spinodal region, the droplet will separate immediately. In addition, if $(\phi_{A}^{ini,dis},\phi_{B}^{ini,dis})$ is outside the miscibility gap but on the same side as the equilibrium point of the continuous phase, the system will evolve towards one single-phase system. Consequently, to reproduce the desired dynamics, the initial compositions of the two phases must necessarily lie on both sides of the spinodal region.
@@
III Simulation setup
In this work, we use the TrioCFD code, an open-source software, developed by the French Alternative Energies and Atomic Energy Commission (CEA) to treat thermal hydraulics problems CEA (2021-2023); Angeli et al. (2015). It has been designed to solve different models including the phase_field model for multiphase flow where coupled Cahn-Hilliard and Navier-Stokes equations are implemented. It is worth noting that in the official version 1.8.4, only binary system with one equation of Cahn-Hilliard can be solved Rasolofomanana et al. (2022b). For the work presented here, we have extended the capabilities of the phase_field application so that it can treat an arbitrary number of components. The numerical results presented here have been shown to be in quantitative agreement with results of simulations using a pseudospectral [^1] No slip boundary conditions were imposed through a diffuse boundary approach. code that has been proven to be very accurate in the context of binary mixtureZanella et al. (2020a, b) and that has been extended to ternary mixtures.
The computational domain consists of a 2D rectangular domain of width $L_{x}=54\times 10^{-3}$ m and $L_{y}=95\times 10^{-3}$ m in which a droplet of initial radius of $R=3.3\times 10^{-3}$ m is placed far enough from the walls to limit edge effects. The wall boundary condition imposes no slip (both components of velocities are zero) at the lower and upper face of the computational domain.

### v2

# Rise and fall of a multicomponent droplet in a surrounding fluid: simulation study of a bumpy path

@@ Abstract
The coupling between mass transfer and hydrodynamic phenomena in two-phase flow is not straightforward due to the different effects that can be encountered. Its description is complex and requires particular efforts, especially in the modelling of the interface between phases. Here, we consider the case of a two components droplet (one miscible and one immiscible in water) released in a 2D rectangular domain filled with water. Mass transfer occurs between the miscible element and the surrounding water, which leads to a density inversion that directly affects the droplet trajectory through buoyancy. We perform simulations using a ternary Cahn-Hilliard model to capture such coupled phenomena. The Boussinesq approximation for a multicomponent system is used to define the density law and an analytical chemical potential is proposed for the thermodynamic landscape. The effect of the mobility parameter on the flow is highlighted, the solutal Marangoni effect is discussed and the results are found in good agreement with the dynamics described from an experimental study of the literature.
Coupled Cahn-Hilliard/Navier-Stokes equations, Multicomponent System, Mass Transfer
@@ I Introduction
Multiphase flows are ubiquitous and can be seen in industrial processes and natural environment. They are difficult to describe and model numerically due to the coupling between the moving free surface and the fluid flow. This leads to complex behaviours that are still a challenge to describe and they are therefore driving new development in modelling. This complexity is augmented when mass transfer between the two phases occurs. This is typically the case in industrial flows and mixing. For instance, this takes place in the pool of molten materials (corium) formed during a severe nuclear accident and affect thermal transfers Zanella et al. (2021); Asmolov et al. (2003); Barrachin and Defoort (2004). It can also occur in microfluidic devices Zhang et al. (2017) and also during the geological sequestration of CO2 either in coal or in deep aquifers De Silva et al. (2015); Zhang and Agarwal (2013). Finally, they are of importance in industrial processes where the dispersion of chemical components in a surrounding fluid is mediated by the motion of droplets Wegener et al. (2014).
@@
In such a situation, mass transfer will occur between the two phases and will affect the fluid flow. This was exemplified in a recent experimental work Rao et al. (2015) inspired by oil spills. Here, through numerical simulations, we show the ability of a ternary phase field model to describe the complex interplay of mass transfer through diffusion and flow equation in the presence of buoyant force. To this purpose, we compare numerical results obtained with a recent experimental work in a qualitative manner. The numerical results also show that the difference between a naive description of the motion and the fully coupled problem can be dramatic and can be well explained by the analysis of the solution of the fully coupled problem. This illustrates the need for fully coupled simulations of such systems. Indeed, simulations methods such as volume of fluids Noh and Woodward (1976) or level set usually rely on the assumption that both phases are immiscible and that no mass transfer occurs through the interface amd S.S. Jain and Dodd (2017). Phase field models based on either the Allen-Cahn Anderson et al. (2000); Folch and Plapp (2005) or the Cahn-Hilliard model Cahn and Hilliard (1958); Kim and Lowengrub (2005); Tóth et al. (2016) can describe mass transfer through the interface. However, this possibility has not been used fully explored and is seen as an adverse effect when modelling multiphase flows of perfectly immiscible fluids Magaletti et al. (2013); Zanella et al. (2020). Here, we show that a ternary Cahn-Hilliard model coupled with the Navier-Stokes equations, using a simple free energy functional is able to describe the complex interplay between fluid flow and diffusive transport and that this modelling effort brings new insights. This is exemplified by numerical simulations of the motion of a multicomponent droplet moving into water while one of its components diffuses out of it leading to changes in the droplet mass density. The numerical results are compared to experimental results presented in Rao et al. (2015).
@@
The paper is organised as follows. First, we briefly describe the phenomena at play in the experiment of interest and discuss qualitatively why a proper description of the coupling between mass transfer and hydrodynamics is needed. Then, we describe the generic form of the model and the simulation setup. We discuss the choice of parameters and propose a convergence study with respect to the interface thickness. Thereafter, we present numerical simulation results that show the ability of the model to describe the interplay between hydrodynamics and mass transfer phenomena before drawing a conclusion and perspectives.
II Description of the physical problem
@@
Indeed, if the droplet composition is sufficiently rich in the light element, it is lighter than water and will move up. Because of the diffusion of the miscible element in water, the amount of A in the droplet decreases while the amount of heavy element B remains constant. As a result, the density of the droplet increases until it reaches the density of pure B which is heavier than water. This implies that the density of the droplet that was initially less than water increases, becomes equal to that of the surrounding water and eventually is larger, resulting in the observed motion.
@@
This explanation is fairly simple, however, if one wants to reproduce this behaviour quantitatively, which may be of interest for instance when modelling flow patterns in corium Zanella et al. (2021), the crude approximation where the diffusive mass transfer is computed considering a droplet in pure water and then the motion of the droplet is obviously incorrect, since it neglects the diffusion layer that builds around the droplet. In the same spirit, considering that the mass transfer is the same as for a quiescent droplet is incorrect, since it neglects the advection that affects the composition around the droplet. Therefore, in order to reproduce quantitatively such motion, one needs to solve the fully coupled problem: the advection-diffusion equation together with the Navier-Stokes equations. This can be done with numerical simulations using the phase field method Kim and Lowengrub (2005); Tóth et al. (2016); Lee et al. (2021); Vasilopoulos (2020). This class of models relies on a diffuse interface that can be tracked implicitly and has been used to model multiphase flows but rarely to model diffusive transport coupled with flow while it has been shown to be efficient Tree et al. (2017).
III Description of the model
@@
The phase field model is a diffuse interface approach. It presents the advantage of being thermodynamically consistent: by construction, it derives from a free energy functional (the thermodynamic landscape). In addition, it is simple to implement since it is based on coupled PDEs. Nevertheless, the thermodynamic landscape and model parameters must be chosen with care to give the right properties. For instance, the diffuse interface thickness is given by a balance between the thermodynamic landscape and the magnitude of a squared gradient term. The accuracy of the model when considering simple multiphase flows has been shown to depend on the diffuse interface thickness and on the mobility of chemical species Magaletti et al. (2013); Zanella et al. (2020). In the more complex case of ternary system, additional interface properties, such as adsorption, have been shown to be difficult to control for thermodynamic landscapes obtained from thermodynamic databases such as the ones constructed by the popular Calphad method Cardon et al. (2016); Rasolofomanana et al. (2022a).
@@
Here, we use the phase field model to study numerically the dynamics of a multicomponent droplet in water. In such approaches, material properties vary continuously from one phase to the other and the interface is defined implicitly as the isosurface of an order parameter. The diffuse interface description was originally proposed by Van Der Waals Rowlinson (1979), it was then popularised by Cahn Cahn and Hilliard (1958) in the context of diffusive mass transfer with a conserved order parameter. It has been then used, with a non conserved order parameter, to describe the motion of free interfaces in a natural way Collins and Levine (1985); Caginalp and Fife (1986) in the context of solid growth where it has become a standard approach Wu (2022); Kadambi et al. (2020). It has been extended thereafter to fluid flow using either a conserved order parameter Anderson et al. (2000) or a non-conserved order parameter Folch and Plapp (2005); Folch et al. (1999). Since the diffusion of chemical species in a ternary system is considered, we use a conserved approach that allows a simple use of ternary system Tóth et al. (2016) and that has already been used in other contexts Tree et al. (2017). It must be noted that in this conserved approach the anticurvature term introduced for non conserved models in Folch et al. (1999) is not needed. In an effort to describe actual systems, there has been recent efforts to couple the principle of such diffuse interface model with free energy functional extracted from thermodynamic databases Cardon et al. (2016); Rasolofomanana et al. (2022a). In this paper, we focus on the description of the interplay between diffusive mass transport and fluid flow. Therefore, we have chosen a free energy functional that is prone to very little interface adsorption.
@@
with $\bm{u}$ the advective velocity field, $\tilde{\mu}_{j}$ the chemical potential of the element $j$, ${G}$ the double-well free energy, the mobility $\mathcal{M}_{i,j}$, the gradient energy $\kappa_{i,j}$ and $\lambda$ an upscaling parameter Cardon et al. (2016); Rasolofomanana et al. (2022a) defined later. The gradient coefficient matrix is chosen simply to be
@@
where $\kappa^{bin}$ is the gradient energy coefficient obtained from a binary system and defined explicitly later. The coefficients of the mobility matrix are taken simply to be equal to:
@@
The choice of $M$, for a given set of parameters is related to the diffusion constant in the bulk phases (either the droplet or the water phase) and is discussed in section V. Here we have chosen to refer to $D_{A}$ the diffusion constant of the miscible element in water.
@@
where $D/Dt=\frac{\partial}{\partial t}+\bm{u}\nabla\cdot$, $P$ is the pressure, $\eta$ is the dynamic viscosity (taken independent of the chemical composition here), $\bm{g}$ is the gravity, the term with $\tilde{\mu}_{i}$ represents the capillary force within the interface as proposed in Kim (2012). Once integrated through the interface at equilibrium, the capillary force leads to the surface tension. $\rho^{\star}$ is the mass density of reference and $\rho$ is the mass density that is function of the order parameters $\phi_{A}$ and $\phi_{B}$ and writes, under the assumption of constant molar volume:
@@
with $\rho^{\star}=$1000 kg/m3, $\beta_{A}=-0.416$ and $\beta_{B}=0.05$, so that the density of pure B is 1050 kg.m-3 and the density of pure A is 584 kg.m-3.
@@
where $\phi_{A}^{eq,cont}$ is the molar fraction of the element A in the continuous phase at thermodynamic equilibrium. With such a choice of thermodynamic landscape, there exists a region where the homogeneous mixture is unstable (the spinodal domain). It corresponds to the domain for which the Hessian matrix of $G$ has one negative eigenvalue and is represented in fig. 1. In this region, small fluctuations of the composition can grow and lead to the formation of domains with different compositions. It should be noted that with this choice of ${G}$, no adsorption at the interface between pure B and pure water occurs when using a diagonal $\bar{\bar{\kappa}}$.
$\begin{array}[]{|c|c||c|c|}\hline\cr\eta&10^{-}3\ \mbox{Pa.s}&\sigma&0.0036\ \mbox{J.m}^{-2}\\ \epsilon&1\mbox{ to }8\ 10^{-4}\mbox{m}&D_{A}&5.4\,10^{-7}\mbox{ to }1.08\,10{-5}\mbox{ m}^{2}.\mbox{s}^{-1}\\ \beta_{A}&-0.416&\beta_{B}&0.05\\ \rho_{A}&584\ \mbox{kg.m}^{-3}&\rho_{B}&1050\ \mbox{kg.m}^{-3}\\ \rho^{\star}&1000\ \mbox{kg.m}^{-3}&&\\ \hline\cr\end{array}$
@@
The initial compositions of the droplet $\phi_{A}^{ini}$ and $\phi_{B}^{ini}$ must satisfy the following conditions. No phase separation should occur during the whole simulation neither in the droplet nor in the bulk phase. Therefore, the initial compositions must be outside the spinodal region and the path between the initial and the equilibrium compositions must not cross it. If ($\phi_{A}^{ini},\phi_{B}^{ini}$) is inside the spinodal region, the droplet will separate immediately. In addition, if $(\phi_{A}^{ini},\phi_{B}^{ini})$ is outside the miscibility gap but on the same side as the equilibrium point of the continuous phase, the system will evolve towards one single-phase system. Consequently, to reproduce the desired dynamics, the initial compositions of the two phases must necessarily lie on both sides of the spinodal region.
@@
IV Simulation setup
In this work, we use an inhouse pseudospectral code that was designed to solve the coupled binary Cahn-Hilliard Navier-Stokes model and has been extended to study ternary Cahn-Hilliard systems for the purpose of this study. The prediction of the code were also tested against the linear stability analysis of the Rayleigh Taylor instability for viscous fluids (at scales where diffusion can be neglected) and very good agreement was reachedZanella et al. (2020, 2021). In addition as described in Supplemental Material [URL to be inserted by the editor] the code was shown to reproduce quantitatively well the initial stages of phase separationCahn (1961) in a ternary system (at scales where flow effect can be neglected when compared to diffusive mass transport).
It is worth noting that the study of this configuration was initiated within the framework of Rasolofomanana’s Ph.D. thesis Rasolofomanana (2023) using the Cahn-Hilliard model Rasolofomanana et al. (2022b) of the TrioCFD code Angeli et al. (2015). This model was extended from binary case to n-ary case for this purpose. The numerical results presented here have been shown to be in good quantitative agreement with results of simulations with TrioCFD code. Only the pseudo-spectral code results are presented in this paper.
The computational domain consists of a 2D rectangular domain of width $L_{x}=54\times 10^{-3}$ m and $L_{y}$, chosen large enough to avoid boundary effects in which a droplet of initial radius of $R=3.5\times 10^{-3}$ m is placed. The wall boundary condition are periodic at top and bottom. In order to avoid a global vertical motion of the fluid the zeroth mode of the buoyant term in the Fourier space is set to zero as in Zanella et al. (2020). In addition a no slip boundary condition at the lateral walls is imposed through the introduction in the Navier-Stokes equation of a dissipative term $-\sigma v$ on a layer of 7 grid points with $\sigma dt=0.5$ which corresponds to an extremely fast relaxation toward 0.